Answer the three parts. If the price charged for a candy bar is p(x) cents, then x thousand candy bars will be sold in a certain city, where p(x) = 1
Answer #1

Given Pra)= 162-17 R(O) = 2-pa = 1620- ② Rix= 162-21 020 x= 162x10 127 810 (11= LO Lo so, mas R1810) = (162x6810)181017 . 10